QUESTION 4 Imran is a dairy farmer and currently has revenues of $34,360 per month. The only costs associated with running the farm are fixed costs of $84,974 per year. As a result of a large shift in demand towards alternative forms of milk in the market Imran serves, he is deciding whether or not to start producing meat instead of milk. He estimates his revenues from producing meat will be $13,403 per month, with no fixed or variable costs. What is Imrans monthly economic profit from producing milk? Answer to the nearest whole number (with no decimal places, $ or- signs, spaces or commas)
Answer #1

Answer 4

Economic profit = Accounting profit - Implicit cost

Accounting profit = Total Revenue - Total cost

If he produces milk His revenue = 34,360 and Total cost = Total fixed cost as Variable cost = 0 => Total cost = 84974/12 = 7081(approx) because annual fixed cost = 84974 => Monthly fixed cost = 84974/12

Hence Accounting profit = 34,360 - 7081 = 27279

Implicit cost is the amount he sacrificed by producing milk. Amount he sacrificed by producing milk is the accounting profit he will earn If instead of producing milk he produces meat.

If he produces meat his Revenue = 13403 and Cost = 0

=> If he produces meat his Accounting profit = 13403 - 0 = 13403

Hence Implicit cost of producing milk = 13403

Thus If he produces milk his Economic profit = Accounting profit - Implicit cost

=> Economic profit = 27279 - 13403 = 13876

Hence, Imran's economic profit from producing milk = 13876